### v2

# Secure numerical simulations using fully homomorphic encryption

Abstract
Data privacy is a significant concern when using numerical simulations for sensitive information such as medical, financial, or engineering data. This issue becomes especially relevant in untrusted environments like public cloud infrastructures. Fully homomorphic encryption (FHE) offers a promising solution for achieving data privacy by enabling secure computations directly on encrypted data. In this paper, aimed at computational scientists, we explore the viability of FHE-based, privacy-preserving numerical simulations of partial differential equations. We begin with an overview of the CKKS scheme, a widely used FHE method for computations with real numbers. Next, we introduce our Julia-based packages OpenFHE.jl and SecureArithmetic.jl, which wrap the OpenFHE C++ library and provide a convenient interface for secure arithmetic operations. We then evaluate the accuracy and performance of key FHE operations in OpenFHE as a baseline for more complex numerical algorithms. Following that, we demonstrate the application of FHE to scientific computing by implementing two finite difference schemes for the linear advection equation. Finally, we discuss potential challenges and solutions for extending secure numerical simulations to other models and methods. Our results show that cryptographically secure numerical simulations are possible, but that careful consideration must be given to the computational overhead and the numerical errors introduced by using FHE.
keywords:
secure numerical simulations , fully homomorphic encryption , CKKS scheme , OpenFHE , Julia programming language
MSC:
[2020] 65-04 , 65M06 , 65Y99
[…]
1 Introduction
Partial differential equations (PDEs) are widely used to model phenomena across nearly every scientific field, from physics and engineering to biology and finance. Since many PDEs cannot be solved analytically, numerical methods are commonly used to approximate their solutions. Beyond strictly scientific settings, PDE-based simulations also play a key role in industry and business applications. In certain cases, the input data and simulation results involve highly sensitive information—–such as personal health records, financial details, or proprietary engineering designs–—that must be carefully protected. For instance, finite element simulations of the aorta used in personalized approaches for treating cardiovascular disease rely on patient-specific data [1, 2]. These datasets must be handled securely to comply with regulations like the General Data Protection Regulation (GDPR) in the European Union or the Health Insurance Portability and Accountability Act (HIPAA) in the United States, posing challenges for scientific research as well as practical implementations [3, 4]. Another example arises in population genetics, where the evolution of gene frequencies has long been modeled by reaction-diffusion equations [5, 6, 7]. Due to nonlinear dynamics and complex boundaries, these models are typically not amenable to analytical solutions and are therefore solved numerically [8, 9, 10]. A practical issue here is working with combined datasets from multiple providers (which do not trust each other), while still protecting the privacy of individual genetic information [11, 12, 13].
To safeguard data, encryption is commonly employed both at rest and during transmission. However, once the data is decrypted for processing, it becomes vulnerable to attacks. The data privacy risks associated with decryption for processing become especially pronounced when data must be processed remotely, as is often the case with public cloud computing platforms. Limited local resources or the need for distributed systems may necessitate offloading computations to external cloud infrastructures, thereby exposing decrypted data to potential attacks. With the rapid expansion of cloud computing [14], ensuring data privacy during computation has emerged as a pressing concern for both practitioners and researchers [15, 16].
A possible solution to this conundrum is the use of what is known as fully homomorphic encryption (FHE). With FHE, it is possible to effect operations directly on the encrypted data. That is, the result of an operation will be same as if the data had been decrypted first, processed, and then encrypted again. Since in case of FHE the data is never available in plaintext, even a fully compromised compute system would not leak any sensitive information. The notion of homomorphic encryption (homomorphisms) was introduced by Rivest, Adleman, and Dertouzos, who in 1978 suggested that such a scheme might be possible [17]. The first viable scheme was proposed much later by Gentry in 2009 [18], when he showed that fully homomorphic encryption is achievable using lattice-based cryptography. Since then, several FHE schemes have been proposed, with the Cheon-Kim-Kim-Song (CKKS) scheme being the most popular option for computations with real numbers [19]. Other commonly used schemes include the Brakerski/Fan–Vercauteren (BFV) [20, 21] and Brakerski–Gentry–Vaikuntanathan (BGV) [22] schemes for computations with integers, and the Ducas–Micciancio (DM/FHEW) [23] and Chillotti–Gama–Georgieva–Izabachene (CGGI/TFHE) [24] schemes for evaluating Boolean and small-integer arithmetic circuits. A good overview of FHE schemes and their applications can be found in [16].
While fully homomorphic encryption schemes represent an active field of research, first applications of FHE schemes can be found in various disciplines, such as privacy-preserving machine learning [25], analysis of encrypted medical or genomic data [26, 27, 28, 13], or for processing data in financial services [29]. One area where FHE has not been applied yet is the field of computational physics. We believe that there are mainly three reasons for this. First, the concept of FHE is still relatively new and not widely known outside of the cryptography community. Second, while there are some excellent open-source libraries available for FHE, they still require a comparatively high level of expertise to use, making it difficult to get started for researchers who are not familiar with the intricacies of FHE. Third, due to its limited scope of available operations, implementing numerical algorithms using FHE is not straightforward and requires a good understanding of FHE.
In this paper, we aim to address these issues by examining fully homomorphic encryption from a scientific computing perspective and introducing it to the computational physics community. We begin with a brief overview of the CKKS scheme for computational scientists, and present two FHE packages we developed in the Julia programming language that allow one to easily experiment with novel FHE applications. We then use these packages to analyze the performance and accuracy of basic FHE operations to get a baseline for more complex algorithms. Finally, we will demonstrate the first prototypes of FHE-secured numerical simulations by implementing and evaluating two well-known numerical methods for solving partial differential equations. We will find that while secure numerical simulations through FHE are indeed possible, the computational overhead is currently significant for many practical scenarios, and the numerical errors introduced by the encryption techniques must be carefully considered when designing applications. All code used in this paper, as well as the data generated during the experiments, is available in our reproducibility repository [30].
The remainder of the paper is structured as follows: In Sec. 2, we give an overview of the CKKS scheme for FHE with real numbers. Moreover, we introduce the open-source C++ library OpenFHE and the Julia packages OpenFHE.jl and SecureArithmetic.jl, which provide a convenient interface for secure arithmetic operations in Julia for experimentation and rapid prototyping. We examine the implementation of the CKKS scheme in OpenFHE in Sec. 3, and analyze the performance and accuracy of basic FHE operations. In Sec. 4, we implement several finite difference schemes to solve the linear advection equation using FHE, and in Sec. 5 we analyze the performance and accuracy of the cryptographically secured numerical solvers. Sec. 6 explores challenges and potential solutions for extending secure numerical simulations to other mathematical models and numerical methods. Finally, we summarize our findings in Sec. 7 and provide an outlook on future work.
2 Fully homomorphic encryption
The fundamental concept of homomorphic encryption is to perform operations on encrypted data without decrypting it first. This is achieved by using a special encryption scheme that allows for the evaluation of arbitrary functions on encrypted data. The outcome of the computation is still encrypted, and needs to be decrypted to obtain the final result. The key property of such an encryption scheme is that the final result is the same as if the encrypted data had first been decrypted, the unencrypted data processed, and the result re-encrypted. Thus, in a sense, a homomorphic encryption scheme is transparent to the application of arbitrary functions to the data. This is illustrated in Fig. 1, where some data $x$ is first homomorphically encrypted, then processed into $f(x)$, and finally decrypted (bottom path). This yields the same result as if the data had been directly processed in the clear (top path).
Figure 1: Comparison of data processing in the clear (top path) to processing it under homomorphic encryption (bottom path). Everything within the hatched area is secure, allowing one to process sensitive information in untrusted environments.
With partially homomorphic encryption, only one type of arithmetic operation can be performed on the encrypted data, i.e., either addition or multiplication. Somewhat homomorphic encryption allows both addition and multiplication, but only supports the evaluation of a subset of computation circuits. Leveled homomorphic encryption extends this to arbitrary computation circuits of bounded (pre-determined) depth. Finally, fully homomorphic encryption (FHE) enables the evaluation of arbitrary functions of unbounded complexity on encrypted data, and is thus the most powerful type of homomorphic encryption. As we will see later, however, there still exist practical limitations to the maximum complexity of functions that can be evaluated using FHE.
In the following, we will first discuss the basics of the CKKS scheme for the fully homomorphic encryption of real numbers and then introduce the software packages we use in this work. Please note that the descriptions of the CKKS scheme and FHE in general are kept at a high level, geared towards computational scientists who are not necessarily experts in cryptography. A good overview of the fundamental principles underlying FHE can also be found in [31].
2.1 Secure approximate arithmetic with the CKKS scheme
The CKKS (Cheon-Kim-Kim-Song) scheme [19] is a fully homomorphic encryption method for approximate arithmetic on real numbers, built upon the BGV scheme for FHE with integers [22]. It obtains its quantum-resistant security from the ring learning with errors (RLWE) problem, which is a specialization of the learning with errors problem [32] for polynomial rings over finite fields [33]. Like other FHE methods, CKKS can be used either as a symmetric or asymmetric encryption scheme. In this manuscript, we will solely focus on its use in the asymmetric encryption context, where there is a public key to encrypt data and a separate private key for decryption.
Fig. 2 outlines the general procedure when using CKKS for FHE. We start with the raw user data $d$, which in case of the CKKS scheme is typically a vector of double-precision floating point numbers (complex numbers are also supported by the CKKS scheme, but they are rarely used in practice). This data is first mapped to a suitable polynomial representation, called the plaintext $p$. The plaintext can then be encrypted using the public key to obtain the ciphertext $c$. The ciphertext $c$ is now ready to be processed by some arbitrary function $f$, which yields another ciphertext $c^{\prime}=f(c)$. Finally, the result can be decrypted using the private key to obtain the plaintext representation $p^{\prime}=f(p)$, and then decoded back to a vector of real numbers $d^{\prime}=f(d)$, which is the desired output.
Figure 2: CKKS overview (following [34]).
The encrypted ciphertext is obtained by adding a random mask generated via RLWE to the plaintext in a controlled manner, effectively hiding the raw data behind a veil of randomness. The computational security of the CKKS scheme, just like any other FHE scheme based on RLWE, relies on the mathematical complexity of solving a noisy high-dimension linear system of equations. This task remains computationally infeasible, even for quantum computers. However, in CKKS the noise also introduces a small error in the computations that is not fully recoverable during the decryption procedure. In this sense, the CKKS scheme is an approximate arithmetic method (which distinguishes it from exact FHE schemes, such as BGV and BFV), and the results of CKKS computations are always subject to some level of approximation error. The magnitude of this error depends on the parameters of CKKS and the number and type of operations performed while evaluating the function $f$.
In terms of arithmetic operations that can be evaluated on the ciphertext, the CKKS scheme supports element-wise addition and multiplication with ciphertexts, plaintexts, or scalar values, as well as rotation operations. The rotation operation performs a circular shift of the data elements in the ciphertext. That is, it moves the data elements by a fixed number of positions while preserving their order, with elements that are shifted out of one end being reintroduced at the opposite end. From these three basic operations, more complex functions can be constructed. Each of these operations further adds a small amount of noise to the data. While this is generally harmless for addition and rotation operations, multiplication adds a non-significant amount of noise, which needs to be kept in check through a process known as rescaling. Due to this, and because of how the CKKS scheme is constructed, the multiplicative depth of computations in CKKS, i.e., the number of consecutive multiplications, is limited to a fixed number of levels. To overcome this limitation, a technique called bootstrapping can be applied after all levels of a ciphertext are used. This bootstrapping procedure resets the available multiplicative depth in the ciphertext, enabling more computations. Since bootstrapping may be used repeatedly, it effectively removes the restrictions on the algorithmic depth of computations in CKKS (for many practical applications where approximation error growth is limited).
Remark 2.1.
With the availability of addition, multiplication, and rotation as arithmetic operations, the CKKS scheme is indeed capable of evaluating arbitrary functions on encrypted data. However, the restriction to just these three basic operations means that, in practice, iterative procedures or polynomial approximations are required to evaluate even moderately complex functions. For example, to determine the multiplicative inverse of a number (e.g., for a division operation), variants of the Newton-Raphson method [35] or Goldschmidt’s algorithm [36] are often used.
In the following sections, we will discuss the CKKS scheme and its main operations in more detail. For clarity, scalar quantities (including polynomials) are written in regular font, while vectors are denoted by bold font. While not strictly necessary to understand the remainder of the paper, these sections provide a useful background for the subsequent sections, especially for readers unfamiliar with FHE methods. Good high-level descriptions of the CKKS scheme can also be found in, e.g., [37, 38, 39].
2.1.1 Plaintext encoding and decoding
The encoding mechanism converts the user data from a representation as a vector of double-precision floating point numbers into a polynomial representation. Polynomials in the CKKS scheme are cyclotomic polynomials $\mathcal{R}=\mathbb{Z}[X]/(X^{N_{R}}+1)$ with $N_{R}$ integer coefficients, where $N_{R}$ is always a power of two and is called the ring dimension. The batch size or capacity, i.e., the number of slots available for packing data into a plaintext, is also a power of two and can be at most as large as half the ring dimension $N_{R}/2$.
For encoding, the input vector is treated as coefficients of a polynomial. This polynomial is first converted to a new representation by evaluating it at complex roots of unity of the $N_{R}$-th cyclotomic polynomial using a procedure similar to an inverse Fast Fourier Transform (FFT). This enables component-wise vector multiplication using FHE. The result of the inverse FFT is then scaled by a large value, the scaling factor (or scaling modulus), and rounded to the nearest integer values to make it compatible with RLWE. Note that at this point, the data is still not not encrypted. Due to the limited precision of (double) floating-point numbers and rounding error, the encoding step already introduces a small error into the data.
The decoding mechanism downscales the integers back to floating-point values and then converts the polynomial back into a data vector by evaluating the polynomial at the complex roots of unity, an operation similar to forward FFT. The noise increase brought about by this operation is typically negligible as the error introduced by the decoding operation is usually smaller than the existing approximation error resulting from prior CKKS operations, such as encryption and computations.
2.1.2 Levels and ciphertext modulus
CKKS is a leveled FHE scheme (with bootstrapping), where a freshly encrypted ciphertext starts with a level $l>0$. With each multiplication of the ciphertext, the level $l$ is decreased by one until $l=0$. Once the level reaches zero, no more multiplications are possible with this ciphertext. Bootstrapping, as discussed in Sec. 2.1.8, alleviates this restriction by resetting the level to a value greater than zero. Addition or rotation operations typically do not consume levels.
Therefore, we usually set FHE parameters using the desired multiplicative depth of the computation. We will denote the maximum multiplicative depth as $l_{\text{max}}$. Based on the multiplicative depth, precision requirements, and security level, two main lattice parameters are configured: ring dimension $N_{R}$ (as described in Section 2.1.1) and ciphertext modulus $Q$.
The ciphertext modulus has the highest value at the beginning, and then gets reduced after each multiplication by the rescaling operation, which is conceptually similar to truncation in floating-point arithmetic, and whose purpose is explained later in Sec. 2.1.6. It is thus convenient to use the notation of $Q_{l}$, where $l$ is the current level. We start with $l=l_{\text{max}}$, i.e., $Q=Q_{l_{\text{max}}}$.
At a more granular level, the ciphertext modulus is a product of $l+1$ factors and is represented as $Q_{l}=q_{0}\cdot q_{1}\cdot q_{2}\dots q_{l-1}\cdot q_{l}$. The rescaling operation reduces the ciphertext modulus from $Q_{l}$ to $Q_{l-1}$. $q_{0}$ is the decryption modulus (often called the first modulus), and for $i=1\dots l$, $q_{i}$ is equal (exactly or approximately) to the scaling factor.
Note that for efficiency, the OpenFHE implementation of CKKS (which we use in this paper, see also Sec. 2.2) uses the Residue Number System (RNS) representation of large numbers, with all values $q_{i}$ being co-prime. For the RNS implementation and a fixed ring dimension $N_{R}$, most of FHE operations scale approximately linearly with the level $l$ [40].
2.1.3 Encryption and decryption
CKKS is a public key encryption scheme with two keys, a public and a private one. The public key $\bm{\mathrm{pk}}$ is used for encryption and can be safely shared, while decryption uses the private key $\bm{\mathrm{sk}}$. The public key is generated using three polynomials: a uniformly random polynomial $a$, a small secret polynomial $s$, and a small error polynomial $e$. The secret key is defined as
$\bm{\mathrm{sk}}=(1,s)$ (1)
and the public key is defined as
$\bm{\mathrm{pk}}\coloneqq(\mathrm{pk}_{0},\mathrm{pk}_{1})=(-a\cdot s+e,a)\pmod{Q_{l}},$ (2)
where $\cdot$ denotes polynomial multiplication. Note that $\mathrm{pk}_{0}\coloneqq-a\cdot s+e$ corresponds to the RLWE problem, which can be informally stated as follows: if $a$ and $\mathrm{pk}_{0}$ are given, it is computationally hard to find $s$ (and $e$). Furthermore, we use $\pmod{Q_{l}}$ to emphasize where arithmetic operations are performed modulo the current ciphertext modulus $Q_{l}$. The error polynomial $e$ is essential for the security of the scheme, as it prevents an attacker from recovering the secret $s$ from the public key $\bm{\mathrm{pk}}$ by solving the corresponding linear systems of equations.
To encrypt the data previously encoded in the plaintext message $\mu$, which itself is a polynomial, three small random polynomials are generated: $u$, $e_{1}$, and $e_{2}$. The public key $\bm{\mathrm{pk}}$ is then used to encrypt the encoded data into a ciphertext $\bm{c}=(c_{0},c_{1})$ as follows:
$c_{0}$ $=\mathrm{pk}_{0}\cdot u+e_{1}+\mu=(-a\cdot s+e)\cdot u+e_{1}+\mu\pmod{Q_{l}},$ (3)
$c_{1}$ $=\mathrm{pk}_{1}\cdot u+e_{2}=a\cdot u+e_{2}\pmod{Q_{l}}.$
As can be seen from Eqn. 3, a ciphertext always consists of a pair of polynomials. Similar to the public key before, the plaintext message is masked by the pseudorandomness coming from the RLWE problem.
[…]
$\tilde{\mu}$ $=\langle\bm{c},\bm{\mathrm{sk}}\rangle=c_{0}+c_{1}\cdot s\pmod{Q_{l}}$ (4)
$=\mu+e\cdot u+e_{1}+e_{2}\cdot s=\mu+v\approx\mu,$
where $v=e\cdot u+e_{1}+e_{2}\cdot s$ is small compared to the plaintext message $\mu$ by construction. Thus, the decryption operation is not exact and adds a small error to the decrypted data.
2.1.4 Addition
Addition of two ciphertexts is done by adding the corresponding polynomials element-wise, i.e.,
$\bm{c}^{\prime}=\bm{c}+\hat{\bm{c}}=(c_{0}+\hat{c}_{0},c_{1}+\hat{c}_{1})\pmod{Q_{l}}.$ (5)
This operation results in an error growth from $e$ to $e+\hat{e}$ in decrypted data, which is still small compared to the data itself. To add a plaintext to a ciphertext, the plaintext is encoded and then extended to the ciphertext space (no additional error is introduced by this operation). Addition of ciphertexts and scalars is done similarly (but no encoding is needed in this case).
2.1.5 Key Switching
Unlike addition, the two other basic arithmetic operations, multiplication and rotation, require executing a special maintenance procedure called key switching. This key switching procedure needs an additional public key, which is often called an evaluation key or key-switching hint/key. We will use the term evaluation key in the paper.
Key switching is necessary because a multiplication or rotation transforms both the encrypted message and the underlying secret key. To transform the resulting ciphertext back to the original secret key, we perform the key switching procedure using a properly generated evaluation key. The evaluation key is essentially an encryption of the transformed secret key under the original secret key, which is done in a way that minimizes the noise increase associated with key switching.
Note that key switching is more computationally expensive in CKKS than the actual multiplication or rotation, and is typically the main bottleneck in FHE applications based on CKKS.
2.1.6 Multiplication
Multiplication of two CKKS ciphertexts $\bm{c}$ and $\hat{\bm{c}}$ is a considerably more complex operation than addition, which is why we will only provide a brief overview here. It requires three steps:
  1. 1.
    Multiplication of the polynomial pairs $\bm{c}=(c_{0},c_{1})$ and $\hat{\bm{c}}=(\hat{c}_{0},\hat{c}_{1})$ to obtain a polynomial triple $\bm{d}=(d_{0},d_{1},d_{2})=(c_{0}\cdot\hat{c}_{0},c_{0}\cdot\hat{c}_{1}+\hat{c}_{0}\cdot c_{1},c_{1}\cdot\hat{c}_{1})$. This is typically called a tensor product.
  2. 2.
    Relinearization of the polynomial triple $d$ to reduce it to polynomial pair $\tilde{\bm{c}}=(\tilde{c}_{0},\tilde{c}_{1})$ again. This step requires a key switching operation.
  3. 3.
    Rescaling of the ciphertext $\tilde{\bm{c}}$ to reduce the message to the same scale as before and to reduce the ciphertext modulus from $Q_{l}$ to $Q_{l-1}$.
As we saw in Sec. 2.1.3, after encryption the first polynomial of the ciphertext $c_{0}$ is linear with respect to the secret key $\bm{\mathrm{sk}}$. After multiplying two ciphertexts using the tensor product, we get a quadratic polynomial in secret key polynomial $s$, i.e., the decryption would be evaluated as $\langle\bm{d},(1,s,s^{2})\rangle=d_{0}+d_{1}\cdot s+d_{2}\cdot s^{2}$. Each subsequent multiplication will increase the degree of $s$ even further; one can think of a multiplication of two ciphertexts as a product of two corresponding decryption polynomials. This implies the size of the ciphertext after a multiplication of two ciphertexts of sizes $i$ and $j$ will grow to $i+j-1$ polynomials. To compress the ciphertext from three to two polynomials (or decrement the size by one in a more general case), the relinearization procedure is needed. In summary, the main goal of relinearization is to maintain the compact representation of ciphertexts.
Another issue is that the multiplication of ciphertext leads to error growth because a small error polynomial is multiplied by encrypted message values. This problem is mitigated by the rescaling technique, which reduces the scaling factor to the one used during encryption, hence truncating the least significant error that appeared as the result of multiplication. Each rescaling reduces the ciphertext modulus $Q_{l}$ by a factor comparable to the scaling factor, thereby decreasing the ciphertext level $l$ and thus the number of multiplications still available.
In leveled FHE, one therefore tries to minimize the required multiplicative depth of an algorithm by evaluating a chain of multiplications using a binary tree approach. For example, instead of computing $\bm{c^{\prime}}=\bm{c}_{1}\cdot\bm{c}_{2}\cdot\bm{c}_{3}\cdot\bm{c}_{4}$ with three consecutive multiplications (and thus requiring three levels), one can first compute $\bm{r}_{1}=\bm{c}_{1}\cdot\bm{c}_{2}$ and $\bm{r}_{2}=\bm{c}_{3}\cdot\bm{c}_{4}$ using one level, and then combine them to $\bm{c^{\prime}}=\bm{r}_{1}\cdot\bm{r}_{2}$ using a second level. In the naïve case, to multiply $n$ ciphertexts, the multiplicative depth of $n-1$ is needed. If the binary tree method is used instead, the multiplicative depth is reduced to approximately $\log_{2}(n)$ (see also [38]).
2.1.7 Rotation
The multiplication and addition operations are element-wise, i.e., they are always performed between corresponding data elements in the user-provided data vectors. These operations are evaluated in a Single-Instruction-Multiple-Data (SIMD) manner, where one homomorphic operation computes many thousands of scalar additions/multiplications at once. Often, we need to perform operations over elements of the same vector or access data elements at particular indices. In these cases, the rotation operation is required, which cyclically shifts encrypted data by some rotation index. When the rotation is evaluated, the underlying secret key gets also rotated. To change the underlying secret key back to the original one (to support further homomorphic computations), one needs to apply the key switching operation. Note that the rotation itself is cheap (effectively it is just reindexing), but the required key switching operation following it is computationally expensive.
An important property of rotation is that it is circular with respect to the batch size/capacity.
2.1.8 Bootstrapping
As described in Sec. 2.1.2 and 2.1.6, the necessity to rescale the ciphertext limits the number of multiplicative operations that can be performed, since no more multiplications are possible once the level $l$ reaches zero. The bootstrapping operation alleviates this restriction by resetting the level to $l>0$. It involves approximating a homomorphic decryption procedure on a ciphertext using an encrypted version of the secret key (this is done indirectly in CKKS), and then re-encrypting it to obtain a refreshed ciphertext. In theory, the number of multiplications that can be performed with a ciphertext is unlimited when using bootstrapping. However, bootstrapping also adds a significant approximation error to the ciphertext and is computationally very expensive. Due to the high complexity of the bootstrapping operation, we will not go into more detail here. In-depth descriptions of various bootstrapping implementations are covered elsewhere [41].
2.2 Software libraries for fully homomorphic encryption
As we have seen, the CKKS scheme is a powerful method for FHE with real numbers. At the same time, it is also an advanced technique that requires a deep understanding of the underlying mathematical methods to be used effectively. It is therefore advisable to use a software library that provides an implementation of the CKKS scheme and other necessary operations. Several such libraries are available, e.g., SEAL [42], HElib [43], and OpenFHE [40].
In this paper, we will focus on the OpenFHE library, which is actively maintained and includes most recent CKKS optimizations [44]. We will also present two Julia packages OpenFHE.jl and SecureArithmetic.jl, which provide a convenient interface to OpenFHE in the Julia programming language. We will now give a brief overview of these tools.
2.2.1 OpenFHE
OpenFHE [40] is an open-source FHE library that offers efficient C++ implementations of all common FHE schemes, including the CKKS scheme with bootstrapping. OpenFHE contains many examples that are especially useful for new users when designing their applications. The OpenFHE library puts an emphasis on usability. For example, the relinearization and rescaling operations after ciphertext multiplication are handled automatically, allowing the library to be used also by non-experts in FHE. This can also be observed in a short code snippet in Listing 2.2.3 (left), where the CKKS internals for multiplication are hidden from the user. Note that some code for initializing the setup was omitted for clarity, e.g., to create the cryptographic context object (named cc in this listing) or the public and private keys. For more details about OpenFHE, please refer to [40].
2.2.2 OpenFHE.jl
The Julia programming language [45] is designed for technical computing, with a simple, expressive syntax and high computational performance. It combines the ease of use of high-level languages like Python with the speed of compiled languages such as C or Fortran, making it ideal for scenarios that require both rapid prototyping and fast execution. Julia comes with its own package manager, which allows users to easily install and manage additional packages, including those written in other languages.
It is this feature that we leverage in our OpenFHE.jl package [46], which provides a Julia wrapper for the OpenFHE library. When installing OpenFHE.jl, pre-built binaries of the OpenFHE library are automatically downloaded without the user having to compile anything locally. The C++ functionality is exposed in Julia via the CxxWrap.jl package [47]. Besides offering a native Julia interface to the OpenFHE library, OpenFHE.jl does not provide any extra functionality. It is intended to be used as a backend for other Julia packages that require FHE capabilities. As can be seen in Listing 2.2.3 (right), the syntax of OpenFHE.jl is very similar to that of OpenFHE, making it easy to port code between the two programming languages.
2.2.3 SecureArithmetic.jl
To make the use of FHE more accessible to non-experts in Julia, we created the SecureArithmetic.jl package [48]. It inherits the functionalities of the OpenFHE.jl package, but provides them consistently through a more convenient Julia syntax. Using the SecureArithmetic.jl package, users can write arithmetic expressions in a common mathematical notation, significantly simplifying the code. Furthermore, SecureArithmetic.jl supports the ability to use the same code for secure and non-secure computations, which is especially useful during the design stage. It allows one to debug FHE algorithms without having to go through the actual encryption and decryption steps, making the execution much faster during the prototyping phase. Listing 2.2.3 shows an example of how the ciphertext multiplication is performed using SecureArithmetic.jl. Compared to OpenFHE/OpenFHE.jl in Listing 2.2.3, SecureArithmetic.jl provides a more high-level interface that is closer to the mathematical notation. However, since it is built on top of OpenFHE.jl and OpenFHE, it is still possible to access the lower-level functions if needed.
{listing}
[!htbp] {minted} [fontsize=, frame=none]cpp std::vector<double> v = 1.0, 2.0, 3.0, 4.0; Plaintext p = cc->MakeCKKSPackedPlaintext(v); auto c = cc->Encrypt(public_key, p); auto c_squared = cc->EvalMult(c, c); Plaintext result; cc->Decrypt(private_key, c_squared, &result); result->SetLength(batch_size); std::cout << "v * v = " << result; // Output: v * v = (1.0, 4.0, 9.0, … {minted} [fontsize=, frame=none]julia v = [1.0, 2.0, 3.0, 4.0] p = MakeCKKSPackedPlaintext(cc, v) c = Encrypt(cc, public_key, p) c_squared = EvalMult(cc, c, c) result = Plaintext() Decrypt(cc, private_key, c_squared, result) SetLength(result, batch_size) println("v * v = ", result) # Output: v * v = (1.0, 4.0, 9.0, … Ciphertext multiplication using OpenFHE in C++ (left) and OpenFHE.jl in Julia (right).
{listing}
[!htbp] {minted}[frame=none]julia v = [1.0, 2.0, 3.0, 4.0] c = encrypt(v, public_key, context) c_squared = c * c result = decrypt(c_squared, private_key) println("v * v = ", result) # Output: v * v = [1.0, 4.0, 9.0, … Ciphertext multiplication using SecureArithmetic.jl in Julia.
3 Accuracy and performance of the CKKS scheme
To provide a baseline for the evaluation of more complex numerical simulations using FHE, in this section we will analyze the accuracy and performance of basic CKKS operations as provided by OpenFHE. In the following, we will first describe the CKKS configuration used throughout this paper and the method we used to measure errors and runtime. We will then present our accuracy and performance analysis for individual FHE operations and discuss the implications of these results for the secure numerical simulations shown later. All code used for the experiments and our numerical results are available in the reproducibility repository [30].
3.1 Experimental setup and measurement methodology
The experiments in this paper were conducted with the OpenFHE library v1.2.0 via the SecureArithmetic.jl package, using a CKKS configuration that provides a 128-bit security level and uses $S_{\text{word}}$ = 64 bits as the native integer size. The security level is determined based on the homomorphic encryption standard [49], and OpenFHE throws an error if any parameters provided violate the security guarantee. All other parameters were chosen on a best-effort basis to balance performance and accuracy. Unless noted otherwise, the configuration used for the OpenFHE setup is as shown in Table 1. While the OpenFHE library has many more user-configurable options, we only focus on those that need to be set explicitly or where we deviate from the default values. For details on the CKKS configuration parameters, please refer to the OpenFHE documentation [40].
Parameter Value
Batch size $2^{5}$$2^{16}$ (as small as possible)
Bootstrapping level budget $[4,4]$
Enabled features ADVANCEDSHE, FHE, KEYSWITCH, LEVELEDSHE, PKE
Size of first modulus $q_{0}$ $60$ bits
Size of scaling modulus $q_{i}$ $59$ bits
Scaling technique FLEXIBLEAUTO
Secret key distribution SPARSE_TERNARY (use UNIFORM_TERNARY for production)
Security level HEStd_128_classic
Table 1: Parameters for the CKKS configuration in OpenFHE. For more details, please refer to the OpenFHE documentation [40].
For the batch size/capacity, we used the smallest power of two that still fits the data. The multiplicative depth available after bootstrapping was chosen to be $l_{\text{refresh}}=15$, with a maximum multiplicative depth of $l_{\text{max}}=33$. The ring dimension is chosen automatically by OpenFHE, which usually corresponds to $N_{R}=2^{17}$ given the remaining parameters. Note that the security standards in the OpenFHE library are being constantly improved, thus in the future a different ring dimension may be supported for these parameters. Furthermore, since the current homomorphic encryption standard [49] does not provide CKKS parameters when using a sparse ternary distribution for the secret key generation, in production settings it is advisable to use the uniform ternary distribution.
The performance and accuracy investigations in this section require applying operations to encrypted and unencrypted data. The main ciphertext is always initialized to
$\left[\sin\left(2\pi\frac{1}{L}\right),\sin\left(2\pi\frac{2}{L}\right),...,\sin\left(2\pi\frac{L}{L}\right)\right],$ (6)
with $L=64$. Scalars are always initialized to $301+\pi/$ 30, and plaintexts to
$\left[1+\frac{\pi}{30},1+\frac{\pi}{30},...,1+\frac{\pi}{30}\right],$ (7)
also with a length of $L=64$. If required, a second ciphertext was created by encrypting the above plaintext. Therefore, the batch size/capacity is always $64$.
Initial tests confirmed that the computational overhead of using OpenFHE through OpenFHE.jl or SecureArithmetic.jl is negligible (below 1%). Therefore, we always used SecureArithmetic.jl to analyze the performance. The experiments were conducted on an AMD Ryzen Threadripper 3990X 64 core processor with 256 GiB RAM using a single thread. All measurements discard the initial just-in-time compilation time of the Julia programming language. We measured the runtime of FHE operations by averaging over five consecutive executions.
Numerical errors are measured by comparing the decrypted result of a secure operation with a corresponding unencrypted operation. In case of encryption, decryption, and bootstrapping operations, the error is calculated relative to doing nothing with the unencrypted data. Secure addition, multiplication, and rotation are compared to regular addition, multiplication, and Julia’s Base.circshift. For the error measurements, we used the $L^{\infty}$ norm, which denotes the maximum absolute element-wise error. Please note that the error measurements should not be interpreted by their exact numbers, but rather by their general size and trend. Different input data or different CKKS configurations may lead to different error values.
3.2 Plaintext encoding and decoding
First, we will consider the accuracy and performance of the encoding and decoding operations. As described in Sec. 2.1.1, the encoding operation is used to transform a vector of real numbers into a plaintext, where it is stored as a polynomial. The decoding is the inverse operation, which converts the polynomial back into a data vector. Unfortunately, the OpenFHE library does not provide a direct way to decode a plaintext (as extra noise is added during decoding for security purposes), thus we can only measure the runtime of the encoding operation, but not the decoding operation, nor the error of either operation. From Fig. 3, we can see that runtime required for encoding a data vector into a plaintext grows linearly with the multiplicative depth $l_{\text{max}}$. While generally the plaintext encoding is independent of $l_{\text{max}}$, by default the OpenFHE library already prepares the plaintext for multiplication with a ciphertext, i.e., transforms the plaintext to ciphertext space. This requires additional expensive operations, which depend on $l_{\text{max}}$, to be performed during encoding. Note that the level used to prepare the plaintext can be configured (changed from its default value) by the user of the library.
Figure 3: Performance evaluation for encoding a vector of real numbers into a plaintext.
3.3 Encryption and decryption
We proceed with analyzing the accuracy and performance of the two most fundamental operations of cryptography: encryption and decryption. Since it is impossible to measure the error of these operations independently, the error was determined after encrypting and immediately decrypting a data vector.
As shown in Fig. 4(a), the error of a single encryption/decryption operation is of $\mathcal{O}(10^{-13})$, and remains approximately there even after more then ten subsequent encryption/decryption pairs. Since the error is approximately in the range of machine precision (for the parameters summarized in Table 1), and usually only a single encryption and decryption are performed, the error is negligible for all practical purposes.
As mentioned in Section 2.1.2, we expect the runtime of ciphertext operations to depend linearly on the multiplicative depth $l_{\text{max}}$. This is confirmed by the results in Fig. 4(b). It is also clear that encryption and decryption take a non-negligible amount of time, already in the order of a second.
(a) Error from encryption/decryption operations.
(b) Time required for encryption/decryption operations.
Figure 4: Error analysis (left) and performance evaluation (right) for encryption and decryption of a ciphertext. For the error investigation, the multiplicative depth was set to $l_{\text{max}}=33$.
Please note that due to the nature of how we measure the accuracy of other operations, the error incurred by one encryption/decryption operation will be included in the error analysis of all other FHE operations in the subsequent sections. Moreover, since it is impossible to measure the error of encoding and decoding operations independently, they are also implicitly included. A more detailed discussion of encryption/decryption errors can be found in [44].
3.4 Addition
Error and runtime measurements for addition operations are presented in Fig. 5. Here, we analyze the addition of two ciphertexts, a ciphertext and a plaintext, and a ciphertext and a scalar. The error of adding a plaintext or a scalar value to a ciphertext is virtually negligible at $\mathcal{O}(10^{-14})$, which is approximately the truncation error for double-precision arithmetic. Adding two ciphertexts incurs an error only slightly larger, on the order of $\mathcal{O}(10^{-13})$ (Fig. 5(a)).
When adding ciphertexts, it is instructive to distinguish between ciphertexts with correlated and uncorrelated errors. To achieve uncorrelated ciphertext errors, one needs to encrypt the second ciphertext summand independently for each operation (or rerandomize the ciphertext by adding an encryption of zero), whereas for correlated ciphertext errors, the same ciphertext is added to itself. As shown in Fig. 5(a), the difference between these two cases is significant. Although the errors initially have the same magnitude, correlated errors increase linearly with the number of additions, as the same errors are repeatedly added. In contrast, uncorrelated errors grow at a rate proportional to the square root of the number of additions, matching the expected behavior based on the Central Limit Theorem.
As with the other operations before, the runtime of the addition operation grows linearly with the multiplicative depth $l_{\text{max}}$. While the addition of two ciphertexts or a ciphertext and a plaintext requires a similar amount of time, adding a scalar to a ciphertext is considerably faster.
(a) Error after addition operations.
(b) Time required for addition operations.
Figure 5: Error analysis (left) and performance evaluation (right) for addition operations between ciphertexts, ciphertexts and plaintexts, and ciphertexts and scalars. For the error investigation, the multiplicative depth was set to $l_{\text{max}}=33$.
3.5 Multiplication
Next, we analyze the accuracy and performance of multiplication operations. As shown in Fig. 6(a), the behavior of the error from multiplying two ciphertexts depends on whether the ciphertexts have correlated or uncorrelated errors. Similar to the findings for ciphertext addition, multiplying ciphertexts with correlated errors leads to a linear increase in overall error, while multiplying ciphertexts with uncorrelated errors results in an error growth proportional to the square root of the number of operations. In both cases, we adjusted the levels of the ciphertexts used for multiplication to maintain consistency across the ciphertexts. In practice, the total error from multiplication in an FHE algorithm will be between the correlated and uncorrelated error curves.
Unlike addition or multiplication of a ciphertext with a plaintext or a scalar, multiplication of two ciphertexts requires a relinearization step, which is especially computationally expensive due to the necessary key switching operation. Therefore, the runtime of ciphertext-ciphertext multiplication increases significantly compared to addition, as can be observed in Fig. 6(b), and is much larger than for multiplication by plaintext or scalar. Furthermore, multiplication of a ciphertext by a plaintext or scalar is much slower than the corresponding addition operations due to the necessary rescaling operation. Generally, one should keep in mind that the runtime for key switching is always significantly larger than the runtime of rescaling, which in turn is significantly larger than the runtimes of the underlying component-wise multiplication or addition.
As before, the runtime grows with increased multiplicative depth $l_{\text{max}}$. Note that for the performance measurements, we evaluated the runtime of the second multiplication of the ciphertext with another factor, since in OpenFHE with the FLEXIBLEAUTO rescaling technique, the rescaling step is not applied to the first multiplication.
(a) Error after multiplication operations.
(b) Time required for multiplication operations.
Figure 6: Error analysis (left) and performance evaluation (right) for multiplication operations between ciphertexts, ciphertexts and plaintexts, and ciphertexts and scalars. For the error investigation, the multiplicative depth was set to $l_{\text{max}}=33$.
3.6 Rotation
The impact of rotation operations on error and execution time is shown in Fig. 7 for different rotation indices. The rotation operation does not result in error growth, thus the error remains at the level of a single encryption-decryption operation. In terms of the runtime, there is still a linear growth with the multiplicative depth $l_{\text{max}}$. Neither the error nor the runtime depend on the rotation index.
(a) Error after rotation operations.
(b) Time required for rotation operations.
Figure 7: Error analysis (left) and performance evaluation (right) for rotation operations by different rotation indices $i\in\{-1,5,-25\}$. For the error investigation, the multiplicative depth was set to $l_{\text{max}}=33$.
3.7 Bootstrapping
Finally, we analyze the accuracy and performance of the bootstrapping procedure. The errors are measured without any arithmetic operations between two subsequent bootstrapping operations. To make the results comparable between different settings for the maximum multiplicative depth $l_{\text{max}}$, we always measure the execution time after dropping the ciphertext to two remaining multiplicative levels $l=2$ before performing the bootstrapping operation.
To improve the accuracy of the bootstrapping operation, in [50] an iterative (or multiprecision) bootstrapping technique was introduced. The idea is to perform multiple bootstrapping operations in sequence when refreshing the ciphertext, each time progressively reducing the error by the precision (in bits) of CKKS bootstrapping. This can significantly increase the bits of precision available for an FHE computation, as demonstrated in, e.g., [51]. The OpenFHE library supports iterative bootstrapping with two consecutive iterations, which requires an experimentally determined precision of a single bootstrapping as an input. In this paper, we use both the standard bootstrapping method with a single iteration and the iterative procedure with two iterations, setting the experimentally determined precision to $19$ bits.
From Fig. 8(a), it becomes clear that bootstrapping has a much higher approximation error than all other CKKS operations, which is consistent with previous findings, e.g., in [41]. The $L^{\infty}$ error at $\mathcal{O}(10^{-6})$ for standard bootstrapping and $\mathcal{O}(10^{-10})$ for iterative bootstrapping is orders of magnitude greater than in all previous experiments (note that adding further bootstrapping iterations is expected to make this error comparable to the error after leveled CKKS operations). When performing multiple bootstrapping procedures in a row, the error seems to remain approximately constant. This may be specific to a relatively small value of the batch size $L=64$ and data points chosen for our experiments (note that in [41] a gradual decline in precision is observed, reaching a plateau for a higher number of bootstrapping operations). Fig. 8(b) further reveals that bootstrapping is also the most time-consuming of all CKKS operations, with a runtime that again grows linearly with the maximum multiplicative depth $l_{\text{max}}$. The runtime for the iterative bootstrapping is approximately twice as large as for the standard bootstrapping. Thus in practice, the increased accuracy of the iterative bootstrapping technique has to be balanced with the additional computational cost. Furthermore, we have to emphasize that the errors and runtimes reported here are highly dependent on the specific CKKS setup and the actual user data, and should thus be only interpreted as indicators and not taken for their absolute values. Unless noted otherwise, in the remainder of the paper we will use the standard bootstrapping method.
(a) Error after bootstrapping operations.
(b) Time required for bootstrapping operations.
Figure 8: Error analysis (left) and performance evaluation (right) for bootstrapping operations. For the error investigation, the multiplicative depth was set to $l_{\text{max}}=33$. The standard bootstrapping procedure uses a single bootstrapping operation, while the iterative procedure performs two subsequent bootstrapping operations.
At this point it should be mentioned that the bootstrapping procedure (as implemented in OpenFHE) requires one remaining multiplicative level per bootstrapping iteration to get started and also consumes multiple levels for the bootstrapping itself. Therefore, after bootstrapping the ciphertext level is restored to $l=l_{\text{refresh}}<l_{\text{max}}$, and the number of levels for continuous operation is $l_{\text{usable}}=l_{\text{refresh}}-1$ for standard and $l_{\text{usable}}=l_{\text{refresh}}-2$ for iterative bootstrapping. Fig. 9 illustrates these values, which are important to consider when designing an FHE application.
Remark 3.1.
Please note that in the OpenFHE library, the value returned by the GetLevel() function does not represent the ciphertext level as it is commonly used in literature and in this paper. Instead, it starts at zero for a fresh ciphertext and increases by one with each multiplication. Also, when using the OpenFHE library with the FLEXIBLEAUTO scaling technique as we are here (see Table 1 above), the rescaling step after a ciphertext multiplication is not performed immediately but only just before the next multiplication. Therefore, the ciphertext level reported by GetLevel() function may be off by one. These implementation-specific peculiarities must be taken into account when implementing algorithms with OpenFHE.
Figure 9: Overview of an exemplary CKKS bootstrapping process with maximum multiplicative depth $l_{\text{max}}=33$. Standard bootstrapping requires one level to get started, and then consumes $18$ levels during the actual process. After bootstrapping, the ciphertext level is restored to $l_{\text{refresh}}=15$ levels, and $l_{\text{usable}}=14$ levels may be used before the next bootstrapping.
We further observed in our experiments that in addition to $l_{\text{max}}$, the error and runtime of bootstrapping are also sensitive to many other parameters, such as the ring dimension $N_{R}$, the batch size/capacity, the ciphertext modulus and scaling factor, the bootstrapping level budget, and the multiplicative depth available after bootstrapping $l_{\text{refresh}}$. It is thus difficult to quantify the accuracy and performance behavior of bootstrapping in general. Moreover, the error also depends on the absolute value of the data in the ciphertext. For algorithms that use bootstrapping, it is thus recommended to rescale the data such that its magnitude is less than one [52]. While it is beyond the scope of this paper to analyze the performance and error implications of all these parameters, we recommend to carefully tailor them to the specific application.
To summarize, of all CKKS operations, bootstrapping is the most detrimental to the accuracy and performance of any FHE application, and should be used as sparingly as possible. However, it is also the ingredient that enables an unlimited number of multiplications in an FHE algorithm (for most of practical applications) and thus facilitates the evaluation of truly arbitrary functions on encrypted data. When the accuracy of the computations is a primary concern, iterative bootstrapping with two iterations can reduce the error by multiple orders of magnitude at the cost of doubling the runtime.
3.8 Overview of accuracy and performance of OpenFHE operations
As noted at the beginning of the accuracy and performance investigation, the exact values of the errors and execution times depend on many factors such as the chosen CKKS parameters, the actual operations performed, the data itself, and on the used compute system. Consequently, all absolute values presented here should be regarded as indicative estimates. We summarize our findings in Table 2, where we analyze the errors and execution times of different FHE operations relative to ciphertext-ciphertext addition. From the table, it is again evident that the most inaccurate and time-consuming FHE operation is bootstrapping. Hence, it is important to use bootstrapping as rarely as possible. On the other hand, the previous results have also shown a linear dependence of the runtime on the maximum multiplicative depth $l_{\text{max}}$ for all FHE operations. Although increasing $l_{\text{max}}$ and the usable levels after bootstrapping $l_{\text{usable}}$ reduces the number of required bootstrapping operations, the runtime of all operations (including bootstrapping itself) increases linearly. Therefore, $l_{\text{max}}$ should be chosen carefully to balance error and runtime, and its optimal value depends on the specific application.
Operation Relative error Relative runtime
Encrypt plaintext $<1$ $\approx 20$
Decrypt ciphertext $<1$ $\approx 30$
Ciphertext + ciphertext $=1$ $=1$
Ciphertext + plaintext $<1$ $\approx 1$
Ciphertext + scalar $<1$ $<1$
Ciphertext * ciphertext $\approx 1$ $\approx 60$
Ciphertext * plaintext $<1$ $\approx 10$
Ciphertext * scalar $<1$ $\approx 6$
Rotate ciphertext $<1$ $\approx 50$
Bootstrap ciphertext (standard) $\approx 30,\!000,\!000$ $\approx 2,\!500$
Bootstrap ciphertext (iterative) $\approx 5,\!000$ $\approx 5,\!000$
Table 2: Error and runtime of OpenFHE operations relative to the addition of two ciphertexts, rounded to the first significant digit. The absolute $L^{\infty}$ error and runtime for ciphertext-ciphertext addition are $\mathcal{O}(10^{-14})$ and $\mathcal{O}(10^{-2})$ s, respectively. Independently encrypted ciphertexts were used for the binary ciphertext operations.
There are some additional measures that can be taken to improve the accuracy and performance of the CKKS scheme. To reduce the error, it is recommended to increase the scaling factor. The iterative bootstrapping method decreases the errors by several orders of magnitude compared to the standard bootstrapping method. Furthermore, the OpenFHE library allows one to set $S_{\text{word}}=128$ bits to internally use larger integers. Most of these measures will, however, increase the runtime of the CKKS operations. To improve the performance, it is advisable to always choose the smallest possible batch size/capacity, since larger batch sizes increase the bootstrapping runtime and reduce the precision. Finally, it is important to find the optimal multiplicative depth $l_{\text{max}}$: a larger one may reduce the frequency of bootstrapping but at the same time increase the size of the ciphertext and, thereby, the execution time of all FHE operations.
4 Setting up secure numerical simulations
The central goal of this work is to demonstrate that fully homomorphic encryption can be used for secure numerical simulations of partial differential equations (PDEs). Since the CKKS scheme only supports three basic arithmetic operations (addition, multiplication, and rotation), we need to recast the numerical methods into a form that can be represented with the available operations. A secondary goal for this manuscript is therefore to start developing the necessary algorithmic building blocks, which will also be useful for other scientists who want to use FHE for their own numerical applications.
In this section, we will begin by introducing the linear advection equation in one and two spatial dimensions, which will serve as the prototypical PDEs for our secure numerical simulations. We will then present two finite difference schemes that we will use to discretize these equations. Next, we will rewrite these schemes in terms of the FHE primitives provided by the CKKS scheme. Finally, we will provide the full algorithm for secure numerical simulations using FHE.
4.1 Linear scalar advection equations
The linear scalar advection equation describes the transport of a scalar field $u$ with a constant speed. In one dimension, it is given by
[…]
where $u=u(t,x)$ is the scalar field, $t$ is time, $x$ is the spatial coordinate, and $a_{x}$ is the advection speed in the $x$-direction. In two spatial dimensions, the linear scalar advection equation reads
[…]
where we now have $u=u(t,x,y)$, $y$ as the second spatial coordinate, and $a_{y}$ as the advection speed in the $y$-direction.
For the simulations, we will use the computational domain $\Omega=[0,1]$ in 1D and $\Omega=[0,1]\times[0,1]$ in 2D. To simplify the FHE implementation, we will consider periodic boundary conditions in all spatial directions. Furthermore, we will restrict our case to positive advection speeds, i.e., $a_{x},a_{y}>0$. A suitable initial condition will be given later for each simulation.
4.2 Finite difference discretization
We use finite difference schemes to approximately solve the linear advection equations in space and time. The computational domain is discretized by a Cartesian mesh with equidistant nodes in each spatial direction as shown in Fig. 10(a) for 1D and in Fig. 10(b) for 2D.
(a) 1D mesh with equidistant nodes on $\Omega=[0,1]$.
(b) 2D mesh with equidistant nodes on $\Omega=[0,1]\times[0,1]$.
Figure 10: Cartesian meshes with equidistant node distributions for the finite-difference discretizations.
The numerical solution $u(t,x)$ is represented in 1D by the values $u_{i}^{n}=u(t_{n},x_{i})$ at the mesh node locations $x_{i}=(i-1)\Delta x$, with $i=1,\dots,N_{x}$, and at time $t_{n}=n\Delta t$, with $n=0,1,2,\dots$ up to the final time. $\Delta x$ is the distance between two neighboring nodes in the $x$-direction and $\Delta t$ is the time step size. In 2D, we have a similar discretization with $u_{ij}^{n}=u(t_{n},x_{i},y_{j})$ at the mesh nodes, where $y_{j}=(j-1)\Delta y$, $j=1,\dots,N_{y}$, and $\Delta y$ being the distance between two neighboring nodes in the $y$-direction.
We use two finite differences schemes to approximately solve the linear advection equation: a first-order upwind scheme and a second-order Lax-Wendroff scheme. The upwind scheme is first-order accurate in space and time. In 1D, the discretization of the linear advection equation is given by
$u_{i}^{n+1}=u_{i}^{n}-\frac{a_{x}\Delta t}{\Delta x}\left(u_{i}^{n}-u_{i-1}^{n}\right),$ (10)
and in 2D by
$u_{ij}^{n+1}=u_{ij}^{n}-\frac{a_{x}\Delta t}{\Delta x}\left(u_{ij}^{n}-u_{i-1j}^{n}\right)-\frac{a_{y}\Delta t}{\Delta y}\left(u_{ij}^{n}-u_{ij-1}^{n}\right).$ (11)
The Lax-Wendroff scheme is second-order accurate in space and time. In 1D, it is given by
$u_{i}^{n+1}=u_{i}^{n}-\frac{a_{x}\Delta t}{2\Delta x}\left(u_{i+1}^{n}-u_{i-1}^{n}\right)+\frac{a_{x}^{2}\Delta t^{2}}{2\Delta x^{2}}\left(u_{i+1}^{n}-2u_{i}^{n}+u_{i-1}^{n}\right)$ (12)
and in 2D by
$\begin{split}u_{ij}^{n+1}=&\left(1-\frac{a_{x}^{2}\Delta t^{2}}{\Delta x^{2}}-\frac{a_{y}^{2}\Delta t^{2}}{\Delta y^{2}}\right)u_{ij}^{n}+\left(\frac{a_{x}^{2}\Delta t^{2}}{2\Delta x^{2}}-\frac{a_{x}\Delta t}{2\Delta x}\right)u_{i+1j}^{n}\\ &+\left(\frac{a_{x}^{2}\Delta t^{2}}{2\Delta x^{2}}+\frac{a_{x}\Delta t}{2\Delta x}\right)u_{i-1j}^{n}+\left(\frac{a_{y}^{2}\Delta t^{2}}{2\Delta y^{2}}-\frac{a_{y}\Delta t}{2\Delta y}\right)u_{ij+1}^{n}\\ &+\left(\frac{a_{y}^{2}\Delta t^{2}}{2\Delta y^{2}}+\frac{a_{y}\Delta t}{2\Delta y}\right)u_{ij-1}^{n}\\ &+\frac{a_{x}a_{y}\Delta t^{2}}{4\Delta x\Delta y}\left(u_{i+1j+1}^{n}-u_{i+1j-1}^{n}-u_{i-1j+1}^{n}+u_{i-1j-1}^{n}\right)\end{split}$ (13)
As stated above, we assume periodic boundary conditions in all spatial directions and that the advection speeds are positive. More details about the employed finite-difference schemes can be found in, e.g., [53].
4.3 Reformulating finite difference schemes for FHE
All FHE operations in the CKKS scheme are vector-based, thus we need to rewrite the finite difference schemes in Eqns. (10)–(13) in a vector-wise manner. Furthermore, we need to limit ourselves to the available operations of element-wise addition, multiplication, and rotation. We will begin with the 1D equations and then proceed to the 2D equations. In the latter case, we will also need to develop a strategy to handle matrix operations with FHE.
In the following, we will again use bold symbols to denote vectors and matrices, e.g., $\bm{u}=(u_{1},u_{2},...,u_{M})^{\intercal}$. Vector/matrix addition and subtraction are written using the usual symbols $+$ and --, respectively. Element-wise vector/matrix multiplication is indicated by the Hadamard product symbol $\odot$. In the spirit of the SecureArithmetic.jl library, we will not distinguish in our notation between unencrypted and encrypted data, since all algorithms work transparently with either.
We represent a generalized rotation operation by $\operatorname{circshift}(\bm{x},k)$, where a vector $\bm{x}$ is cyclically shifted by $k$ positions. If $k$ is positive, the shift moves elements forward; otherwise, it moves them backward. For example,
$\operatorname{circshift}\left(\begin{pmatrix}a\\ b\\ c\end{pmatrix}\!\!,1\right)=\begin{pmatrix}c\\ a\\ b\end{pmatrix}.$ (14)
For algorithms in two dimensions, we need to extend the circular shift operation to matrices. In case of a matrix $\bm{A}\in{\mathbb{R}}^{n\times m}$, $\operatorname{circshift}(\bm{A},k,l)$ denotes a cyclic shift by $k$ rows and $l$ columns. For example,
$\operatorname{circshift}\left(\begin{pmatrix}a&b&c\\ d&e&f\\ g&h&i\end{pmatrix}\!\!,1,2\right)=\begin{pmatrix}h&i&g\\ b&c&a\\ e&f&d\end{pmatrix}.$ (15)
Depending on the context, the “$\operatorname{circshift}$” operator may represent either the rotation of unencrypted or encrypted vectors and matrices. When we specifically refer to the CKKS rotation operation, we will use the term “$\operatorname{rotate}$”. For our goal to rewrite the finite difference schemes in terms of operations supported by FHE, we thus have addition/subtraction, multiplication, and $\operatorname{circshift}$ at our disposal.
4.3.1 Rewriting the finite difference methods in 1D
With the definitions above, we can rewrite the 1D first-order upwind finite difference scheme from Eqn. (10) with periodic boundary conditions as
$\bm{u}^{n+1}=\bm{u}^{n}-\frac{a_{x}\Delta t}{\Delta x}\Bigl{(}\bm{u}^{n}-\operatorname{circshift}(\bm{u}^{n},1)\Bigr{)}.$ (16)
Similarly, we can rewrite the 1D Lax-Wendroff scheme from Eqn. (12) with periodic boundary conditions as
$\begin{split}\bm{u}^{n+1}=\bm{u}^{n}&-\frac{a_{x}\Delta t}{2\Delta x}\Bigl{(}\operatorname{circshift}(\bm{u}^{n},-1)-\operatorname{circshift}(\bm{u}^{n},1)\Bigr{)}\\ &+\frac{a_{x}^{2}\Delta t^{2}}{2\Delta x^{2}}\Bigl{(}\operatorname{circshift}(\bm{u}^{n},-1)-2\bm{u}^{n}+\operatorname{circshift}(\bm{u}^{n},1)\Bigr{)}\end{split}$ (17)
Recall from Sec. 2.1.1 that in the CKKS scheme, vectors of real numbers are encoded into plaintexts with a certain batch size/capacity, which is at most half the ring dimension $N_{R}$. If the batch size coincides with the length of the solution vector $\bm{u}$, we can directly use the CKKS rotation operation for our $\operatorname{circshift}$ operation. In general, however, we need a way to handle the rotation of vectors where the length of the user-provided vector is smaller than the capacity. This is especially relevant for practical applications, since the batch size is always a power of two.
Therefore, we need to augment the $\operatorname{circshift}$ operation to support the case where the length of the actual data vector is less than the capacity of the ciphertext. We present an algorithm for this in Alg. 1. The algorithm performs two rotations of the input ciphertext $\bm{x}$, one backward and one forward, and then uses multiplication with appropriate masking vectors to combine the elements from both rotations. In the algorithm, we designate the CKKS rotation operation as “$\operatorname{rotate}$”. Since the OpenFHE library uses a different sign convention for the rotation index than we use for the $\operatorname{circshift}$ operation (which follows the sign convention of the corresponding function in the Julia base library, Base.circshift), we will need to adjust the sign of the shift in the $\operatorname{rotate}$ operation. Note that for the input vectors in Alg. 1, we consider the length to be equal to the length of the actual user data, while the capacity is equal to the batch size.
Algorithm 1 Circular shift for secure vectors.
[…]
2:
3:$\bm{u}\leftarrow\operatorname{rotate}(\bm{x},-k)$
[…]
6:$f_{1}\leftarrow 1$
7:$l_{1}\leftarrow\text{length}(x)+k$
8:$f_{2}\leftarrow 1+\text{length}(x)+k$
9:$l_{2}\leftarrow\text{length}(x)$Else
10:$f_{1}\leftarrow 1+k$
11:$l_{1}\leftarrow\text{length}(x)$
$f_{2}\leftarrow 1$
13:$l_{2}\leftarrow k$EndIf
14:
15:$\bm{m}_{1}\leftarrow(0,\dots,0)^{\intercal}$triangleright Create masking vectors as plaintexts
[…]
17:$\bm{m}_{2}\leftarrow(0,\dots,0)^{\intercal}$
[…]
19:
20:$\bar{\bm{u}}\leftarrow\bm{u}\odot\bm{m}_{1}$$\triangleright$ Apply masks by multiplying ciphertexts with plaintexts
21:$\bar{\bm{v}}\leftarrow\bm{v}\odot\bm{m}_{2}$
22:
[…]
With this implementation for the $\operatorname{circshift}$ operation, we now support all operations to compute the solution at the new time step in Eqns. (16) and (17). Note that $\operatorname{circshift}$ consumes one multiplicative level if the length of the ciphertext is less than its capacity, and zero levels otherwise. Therefore, a single evaluation of Eqn. (16) or Eqn. (17) uses two multiplicative levels if the length of the ciphertext $\bm{u}$ is less than the capacity, and one level otherwise. Alg. 1 is also the basis for the $\operatorname{circshift}$ implementation in the SecureArithmetic.jl package.
4.3.2 Matrix arithmetic with FHE ciphertexts
To reformulate the 2D finite difference schemes in matrix-vector form, we first need to establish a means of representing data as a matrix in CKKS. Unfortunately, currently no FHE scheme natively supports matrix arithmetic, thus we need to find a vector-based matrix representation and define arithmetic operations on it.
The simplest solution would be to put every row of a matrix in a separate ciphertext. Unfortunately, this approach quickly becomes inordinately inefficient for problems with a user data length that would otherwise fit into a single ciphertext. Consider, for example, a 2D domain discretized by $40\times 40$ nodes. Since the $40\times 40$ matrix has only $1600$ elements, it could still be held in a single ciphertext with ring dimension $N_{R}=2^{17}$. However, we can also store the solution matrix as $40$ vectors with a length of $40$, each in a separate ciphertext. In the latter case, each FHE operation would require approximately $40$ times longer, since it would have to be performed for each ciphertext separately. Furthermore, due to security requirements and other algorithmic constraints, such a storage strategy still results in ciphertexts with a ring dimension of $2^{17}$. Thus, the size of the matrix in memory will be around $2.5$ GB for a multiplicative depth of $l_{\text{max}}=31$, as opposed to a single ciphertext with $64$ MB. Obviously, we need to find another vector-based matrix representation.
Therefore, we store a matrix by fusing its elements column by column and storing it in a single ciphertext. This is akin to programming languages that store matrix data linearly in memory in a column-major order. The procedure is visualized in Fig. 11. Since the batch size/capacity of CKKS ciphertexts is always a power of two, in this example the last four slots in the ciphertext are unused and of indeterminate value.
Figure 11: Storing a $4\times 3$ matrix (left) in a one-dimensional ciphertext with capacity $16$ (right) by fusing its columns in column-major order. The last four slots of the ciphertext remain unused.
The good news is that ciphertext addition and multiplication in CKKS are element-wise, and thus element-wise matrix addition and multiplication are supported out of the box. However, the $\operatorname{circshift}$ operation does not extend directly to the matrix, since the underlying CKKS $\operatorname{rotate}$ operation performs only one-dimensional rotations.
We therefore need to extend Alg. 1 to two dimensions. One possible approach is given in Alg. 2. Similar to the 1D algorithm, it works through a series of rotations and appropriate masking vectors, which are then additively combined for the final result. With the “size” function we obtain the number of rows and columns of the matrix. Note that the two-dimensional $\operatorname{circshift}$ operation consumes between zero and two multiplicative levels, depending on whether a mask is required and if the underlying ciphertext holding the matrix data has a length that is equal or less than its capacity. An overview of the required number of multiplications is given in Table 3. Note that Alg. 2 is also the basis for the $\operatorname{circshift}$ implementation for ciphertexts representing matrix data in the SecureArithmetic.jl package.
Algorithm 2 Circular shift for secure matrices.
[…]
2:return $\operatorname{circshift}(\bm{x},ln)$EndIf
[…]
4:$f_{1}\leftarrow 1$
5:$l_{1}\leftarrow n-k$
6:$f_{2}\leftarrow 1+n-k$
7:$l_{2}\leftarrow n$Else
8:$f_{1}\leftarrow 1-k$
9:$l_{1}\leftarrow n$
10:$f_{2}\leftarrow 1$
11:$l_{2}\leftarrow-k$EndIf
12:
13:$\bm{m}_{\text{tmp},1}\leftarrow(0,0,\dots,0)^{\intercal}\in\mathbb{R}^{n}$triangleright Create masking vectors for one column as plaintexts
[…]
15:$\bm{m}_{\text{tmp},2}\leftarrow(0,0,\dots,0)^{\intercal}\in\mathbb{R}^{n}$
[…]
17:$\bm{m}_{1}\leftarrow$repeat$(\bm{m}_{\text{tmp},1},m)\in\mathbb{R}^{nm}$triangleright Repeat the mask for each column
18:$\bm{m}_{2}\leftarrow$repeat$(\bm{m}_{\text{tmp},2},m)\in\mathbb{R}^{nm}$
19:
20:$\bar{\bm{u}}\leftarrow\bm{x}\odot\bm{m}_{1}$$\triangleright$ Apply masks by multiplying ciphertexts with plaintexts
21:$\bar{\bm{v}}\leftarrow\bm{x}\odot\bm{m}_{2}$
22:
23:$s_{1}\leftarrow ln+k$triangleright Compute one-dimensional shift indices
[…]
26:return $\operatorname{rotate}(\bar{\bm{u}},-s_{1})+\operatorname{rotate}(\bar{\bm{v}},-s_{2})$Else
[…]
$k$ $l$ Levels
$=0$ $=0$ 00
$=0$ $\neq 0$ $1$
$\neq 0$ $=0$ $1$
$\neq 0$ $\neq 0$ $2$
(a) Length $\neq$ capacity.
$k$ $l$ Levels
$=0$ $=0$ 00
$=0$ $\neq 0$ 00
$\neq 0$ $=0$ $1$
$\neq 0$ $\neq 0$ $1$
(b) Length $=$ capacity.
Table 3: Multiplicative levels required by two-dimensional $\operatorname{circshift}$ for rotation by $k$ rows and $l$ columns.
4.3.3 Rewriting the finite difference methods in 2D
With the algorithmic building blocks of Sec. 4.3.2, we can finally rewrite the 2D upwind scheme from Eqn. (11) with periodic boundary conditions in matrix-vector formulation as
$\begin{split}\bm{u}^{n+1}=\bm{u}^{n}&-\frac{a_{x}\Delta t}{\Delta x}\Bigl{(}\bm{u}^{n}-\operatorname{circshift}\bigl{(}\bm{u}^{n},1,0\bigr{)}\Bigr{)}\\ &-\frac{a_{y}\Delta t}{\Delta y}\Bigl{(}\bm{u}^{n}-\operatorname{circshift}\bigl{(}\bm{u}^{n},0,1\bigr{)}\Bigr{)}\end{split}$ (18)
Similarly, we can rewrite the 2D Lax-Wendroff scheme from Eqn. (13) with periodic boundary conditions as
$\begin{split}\bm{u}^{n+1}=\Bigl{(}1-\frac{a_{x}^{2}\Delta t^{2}}{\Delta x^{2}}-\frac{a_{y}^{2}\Delta t^{2}}{\Delta y^{2}}\Bigr{)}\bm{u}^{n}&+\Bigl{(}\frac{a_{x}^{2}\Delta t^{2}}{2\Delta x^{2}}-\frac{a_{x}\Delta t}{2\Delta x}\Bigr{)}\operatorname{circshift}\bigl{(}\bm{u}^{n},-1,0\bigr{)}\\ &+\Bigl{(}\frac{a_{x}^{2}\Delta t^{2}}{2\Delta x^{2}}+\frac{a_{x}\Delta t}{2\Delta x}\Bigr{)}\operatorname{circshift}\bigl{(}\bm{u}^{n},1,0\bigr{)}\\ &+\Bigl{(}\frac{a_{y}^{2}\Delta t^{2}}{2\Delta y^{2}}-\frac{a_{y}\Delta t}{2\Delta y}\Bigr{)}\operatorname{circshift}\bigl{(}\bm{u}^{n},0,-1\bigr{)}\\ &+\Bigl{(}\frac{a_{y}^{2}\Delta t^{2}}{2\Delta y^{2}}+\frac{a_{y}\Delta t}{2\Delta y}\Bigr{)}\operatorname{circshift}\bigl{(}\bm{u}^{n},0,1\bigr{)}\\ &+\frac{a_{x}a_{y}\Delta t^{2}}{4\Delta x\Delta y}\Bigl{(}\operatorname{circshift}\bigl{(}\bm{u}^{n},-1,-1\bigr{)}-\operatorname{circshift}\bigl{(}\bm{u}^{n},-1,1\bigr{)}\\ &\qquad\qquad\quad\,\,\,-\operatorname{circshift}\bigl{(}\bm{u}^{n},1,-1\bigr{)}+\operatorname{circshift}\bigl{(}\bm{u}^{n},1,1\bigr{)}\Bigr{)}\end{split}$ (19)
Note that $\bm{u}^{n+1}$ is a matrix now, with coefficients $u_{ij}$ at each node location. Consequently, we also need to use the matrix representation of real data in a ciphertext and the matrix version of the $\operatorname{circshift}$ operation, which were introduced in the previous section.
4.4 Full algorithm for FHE-secured simulations
All introduced numerical schemes in Eqns. (16)–(19) are one-step methods. Alg. 3 represents the general structure of how all such numerical methods can be implemented with FHE. Here, it is assumed that the initial solution $\bm{u}^{0}$ is already encrypted. The algorithm iterates over the time span $t_{0}$ to $t_{\text{end}}$ with a fixed time step $\Delta t$. In each iteration, we check if during the next iteration (requiring $l_{\text{step}}$ multiplicative levels), the ciphertext level would would fall below one (or two, if iterative bootstrapping is used). If this is the case, the solution is bootstrapped to increase the ciphertext level again. The actual computation of the solution at the next time step is then performed in line 6. The algorithm is terminated when the final time $t_{\text{end}}$ is reached.
Algorithm 3 FHE algorithm for secure numerical simulations.
1:$\bm{u}\leftarrow\bm{u^{0}}$triangleright Initialize solution
2:$t\leftarrow t_{0}$triangleright Initialize time
[…]
4:$\bm{u}\leftarrow$bootstrap$(\bm{u})$EndIf
5:
6:$\bm{u}^{n+1}\leftarrow f(\bm{u})$triangleright Compute solution at next time step based on Eqns. (16)–(19)
7:$\bm{u}\leftarrow\bm{u}^{n+1}$
8:
[…]
Compared to a classical numerical simulation implementation, the main difference is the check for the multiplicative depth and the execution of the bootstrapping operation if required. One should keep in mind that since branching on encrypted ciphertext values is impossible by design, the condition for exiting a loop cannot be cryptographically secured. In our implementation we chose to keep the time and step size a plaintext value. To also hide this information from a potential adversary, one could prescribe a fixed number of iterations instead, after which the result is returned. This is also the only option if, in a modified algorithm, the time step were to be computed dynamically from the encrypted solution data. In addition, we made the decision to further keep some the of the numerical setup as plaintext, such as the advection speeds or the spatial step size. This is due to the fact that division is not natively supported by FHE, and thus factors such as $a_{x}\Delta t/\Delta x$ need to be computed a priori. However, the actual solution data is always kept in encrypted form.
The accuracy and performance of cryptographically secure numerical simulations depends significantly on the performed FHE operations. For Alg. 3 and the numerical schemes introduced in the previous sections, we conducted a static analysis of the number of FHE operations (add, multiply, rotate, and bootstrap) that are executed in each iteration. Our findings are summarized in Table 4.
length $\neq$ capacity
Upwind Lax-Wendroff
1D 2D 1D 2D
$3$ $6$ $4$ $24$
$3$ $6$ $7$ $38$
$2$ $4$ $4$ $24$
$2/l_{\text{usable}}$ $2/l_{\text{usable}}$ $2/l_{\text{usable}}$ $3/l_{\text{usable}}$
CKKS
operation
Add
Multiply
Rotate
Bootstrap
length $=$ capacity
Upwind Lax-Wendroff
1D 2D 1D 2D
$2$ $5$ $2$ $14$
$1$ $4$ $3$ $18$
$1$ $3$ $2$ $14$
$1/l_{\text{usable}}$ $2/l_{\text{usable}}$ $1/l_{\text{usable}}$ $2/l_{\text{usable}}$
Table 4: Number of FHE operations required for a single time step for different numerical schemes.
As expected, the two-dimensional schemes require more FHE operations than their one-dimensional counterparts. Similarly, the higher-order Lax-Wendroff scheme requires more operations than the first-order upwind scheme. The number of operations also depends on the length of the user data: if it is equal to the ciphertext capacity, some of the circular shifts reduce to simple CKKS rotations, which necessitate fewer operations than a full $\operatorname{circshift}$ operation. In addition, using a data size less than the ciphertext capacity increases the multiplicative depth due to additional masking operations.
While the higher-order Lax-Wendroff scheme is more accurate in theory, it also uses more FHE operations per iteration, which increases both runtime and the error introduced by the CKKS scheme. Therefore, these requirements need to be balanced against each other in practice, and they likely depend on the specific application. A more detailed error and performance analysis for our particular setup will be presented in the next section.
5 Numerical results and performance of secure numerical simulations
With all ingredients for FHE-compatible algorithms in place, we can now proceed to conducting secure numerical simulations. Our investigation aims to compare the accuracy and performance of the cryptographically secure simulations with their classical counterparts. For the CKKS setup in OpenFHE, we generally use the same parameters as in Sec. 3.1, but with the available depth after bootstrapping being set to $l_{\text{refresh}}=25$. The ring dimension remains at $N_{R}=2^{17}$. All simulations were conducted using SecureArithmetic.jl with a single thread unless noted otherwise. The code for the simulations and the results are available in our reproducibility repository [30].
For the numerical simulations presented here, the advection speed of the linear advection equations is set to $a_{x}=a_{y}=1$. The initial condition is given by $u_{0}(x)=\sin(2\pi x)$ in 1D and $u_{0}(x,y)=\sin(2\pi x)\sin(2\pi y)$ in 2D, and periodic boundary conditions are imposed. Unless noted otherwise, we will use $N=N_{x}=N_{y}=64$ equidistant nodes in each spatial direction to discretize the computational domain, and a simulation time span of $t\in[0,1]$. The time step size $\Delta t$ is determined from the CFL condition with a CFL number of $0.5$.
In the following, we will first conduct convergence tests to verify that the FHE-secured numerical simulations do not break the convergence properties of the numerical schemes, and present the numerical results of a full simulation. We will then discuss the accuracy and performance of the secure simulations and compare them to their unencrypted counterparts. Finally, we will briefly explore the potential for speeding up the simulations by leveraging the multi-threading capabilities of the OpenFHE library.
5.1 Convergence test
We begin by performing a convergence study to verify the implementation of the numerical methods and to make sure that the convergence properties of the numerical schemes are not affected by the FHE encryption. We will use the simulation setup as described, but vary the number of nodes in each direction from $N=32$ to $N=256$ and reduce simulation time span to $t\in[0,0.5]$. To speed up the convergence tests, we ran the simulations in parallel with eight threads (for more details on parallelizing OpenFHE with OpenMP, see Sec. 5.4 below).
We use the discrete $L^{2}$ norm as a measure of the error between the numerical solution $u$ and the exact solution $u_{\text{exact}}$, where $u_{\text{exact},i}^{n}=u_{0}(x_{i}-a_{x}t_{n})$ in 1D and $u_{\text{exact},ij}^{n}=u_{0}(x_{i}-a_{x}t_{n},y_{j}-a_{y}t_{n})$ in 2D. The $L^{2}$ error for $N$ nodes per direction is then calculated by
[…]
From the $L^{2}$ error, we then compute the experimental order of convergence (EOC) at each step by
$\text{EOC}=\log_{2}(e_{N}/e_{2N})$ (21)
Table 5 contains the results of the convergence study. The EOC values are close to the expected values of one for the first-order upwind scheme and two for the second-order Lax-Wendroff scheme. This confirms the correctness of the implementation and shows that the FHE encryption does not break the convergence properties of the numerical schemes.
1D, upwind $N$ $L^{2}$ error EOC 32 1.01e-01 - 64 5.25e-02 0.95 128 2.67e-02 0.97 256 1.35e-02 0.99 1D, Lax-Wendroff $N$ $L^{2}$ error EOC 32 1.07e-02 - 64 2.67e-03 2.00 128 6.69e-04 2.00 256 1.67e-04 2.00 2D, upwind $N$ $L^{2}$ error EOC 32 1.88e-01 - 64 1.07e-01 0.82 128 5.69e-02 0.90 256 2.94e-02 0.95 2D, Lax-Wendroff $N$ $L^{2}$ error EOC 32 1.07e-02 - 64 2.68e-03 2.00 128 6.69e-04 2.00 256 1.67e-04 2.00
Table 5: Convergence test results for the 1D and 2D discretizations of the linear scalar advection equation with the first-order upwind scheme and the second-order Lax-Wendroff scheme.
5.2 Results for a full simulation
Fig. 12 shows the results for a full simulation of the linear scalar advection equation in 1D at $t=1$, i.e., after one full period of this periodic problem setup. The results of the upwind scheme and the Lax-Wendroff scheme are compared to the exact solution. As expected, the first-order upwind scheme is much more dissipative than the second-order Lax-Wendroff scheme, with the latter being visually indistinguishable from the exact solution.
Figure 12: Secure simulation results of the linear scalar advection equation with the first-order upwind and the second-order Lax-Wendroff schemes in 1D at $t=1$.
In Fig. 13, we present the simulation results for the linear scalar advection in 2D, also at $t=1$. Again, the first-order upwind scheme displays significant dissipative effects. We omitted the exact solution, since it is again indiscernible from the Lax-Wendroff scheme.
(a) First-order upwind scheme.
(b) Second-order Lax-Wendroff scheme.
Figure 13: Secure simulation results of the linear scalar advection equation in 2D at $t=1$.
5.3 Accuracy and performance of encrypted vs. unencrypted simulations
With the convergence properties of the numerical solvers verified, we proceed with comparing the accuracy of the results and the performance of the secure algorithms with their unencrypted counterparts. We continue to use the same simulation setup as given above. Since all algorithms are implemented by using SecureArithmetic.jl, we can directly compare the results by running the simulations once with the OpenFHE/OpenFHE.jl backend and once with an unencrypted backend that internally uses plain Julia vectors/matrices. This allows us to measure the error and runtime behavior of the encrypted simulations relative to the unencrypted on otherwise identical implementations.
In Fig. 14 we analyze the accuracy of the cryptographically secure numerical simulations compared to the unencrypted simulations. Note that unlike the convergence test, where we used the exact solution as a reference, we now compare the results of the secure simulations to the results of the unencrypted simulations. This allows us to separate the error of the numerical discretization and instead only consider the error introduced by the FHE encryption, which we compute as the $L^{\infty}$ norm of the difference between the secure and unencrypted solutions at each time step. In both the 1D and 2D cases, we can see that the error for the Lax-Wendroff scheme is higher than for the upwind scheme. This is expected, since the Lax-Wendroff scheme requires more FHE operations, which in turn increases the error introduced by the CKKS scheme. As we saw above, however, overall the Lax-Wendroff scheme is still much more accurate than the upwind scheme, even with the additional error introduced by the encrypted operations. Furthermore, the relative $L^{\infty}$ error at $\mathcal{O}(10^{-6})$ is still several orders of magnitude lower then the absolute numerical error of the unencrypted simulations at $\mathcal{O}(10^{-3})$. Yet, it is also clear that with FHE, less than single-precision floating point accuracy is achieved, which may be a limiting factor for some applications.
(a) 1D.
(b) 2D.
Figure 14: Errors introduced by the CKKS scheme for the numerical simulations with different discretization methods.
It is also very interesting to see the large effect the bootstrapping operation has on the accuracy. Until the first bootstrapping operation at around 40 time steps in the 1D simulations and at around 25 time steps in the 2D simulations, the error is very low at $\mathcal{O}(10^{-13})$. After the bootstrapping operation, the error jumps by several orders of magnitude to $\mathcal{O}(10^{-6})$. We also observe that the error seems to gradually decrease again after the bootstrapping operation, which is especially notable for the 2D upwind solution. The observation that the precision of CKKS with bootstrapping improves over time is a relatively common phenomenon in privacy-preserving machine learning applications of CKKS. For instance, Han et al. show that the CKKS error in encrypted logistic regression training is higher in initial iterations, but then the encrypted solution gradually converges to the plaintext result due to the convergence of the underlying (logistic regression training) procedure (see Section 5.2 of [54]). We expect similar behavior in converging finite difference schemes, and consider the observed precision improvement after bootstrapping as an indicator of this more general phenomenon.
Next, we look at the runtime performance of the secure numerical simulations. In Table 6(a) we show the runtime per step for the different schemes, which in case of the secure simulations was measured between the first and the second bootstrapping operation, including bootstrapping, and then averaged over the number of time steps. From the table, we can clearly see that the secure simulations are significantly slower than the unencrypted simulations, by approximately six orders of magnitude. Considering the runtime measurements of the individual CKKS operations in Sec. 5.3, this is not surprising. Furthermore, the 2D simulations are more expensive than the 1D simulations, and the Lax-Wendroff scheme has a higher runtime than the upwind scheme, which also matches expectations. It is interesting to note that the runtime of the encrypted simulations only differs by a factor of two to four between the 1D and 2D simulations, while the unencrypted 2D simulations are around ten times more expensive than their 1D counterparts. Considering that the FHE operation count for the 2D simulations is much higher than for the 1D simulations, as presented in Table 4, this is a strong indication of the impact the bootstrapping operation has on the runtime of the secure simulations.
OpenFHE Unencrypted
1D Upwind 4.67 s 6.85e-7 s
Lax-Wendroff 5.28 s 1.21e-6 s
2D Upwind 10.5 s 2.41e-5 s
Lax-Wendroff 20.8 s 6.92e-5 s
(a) Runtime per time step.
OpenFHE Unencrypted
1D Upwind 93.3 s 4.44e-6 s
Lax-Wendroff 101.0 s 5.89e-6 s
2D Upwind 170.1 s 9.73e-4 s
Lax-Wendroff 220.8 s 1.09e-3 s
(b) Time for initialization.
Table 6: Runtime per time step (left) and initial configuration time (right) for the 1D and 2D simulations with the secure and insecure backends.
A brief analysis of the time required for the initialization of the CKKS scheme is given in Table 6(b). Not surprisingly, the initial setup of the FHE operations is orders of magnitude slower than the initialization of the unencrypted data structures. Even though ciphertexts with the same ring dimensions are used, the 2D times are much higher than for the 1D simulations, owing to the fact that they have a significantly larger batch size/capacity. In addition, the 2D $\operatorname{circshift}$ operation requires a larger number of index shifts, which need to be set up during the initialization phase. Similarly, the Lax-Wendroff scheme requires rotations in more directions than the upwind scheme, which also increases the initialization time.
Next to the average runtime per step, it is instructive to look at the runtime over the number of time steps (Fig. 15). Especially for the 1D simulations in Fig. 15(a), we can see that the run time per step decreases with the number of steps, until there is a sudden and significant increase in runtime. These jumps can be attributed to the bootstrapping operation, matching the first bootstrapping after around 40 steps and then every approximately 25 steps thereafter. This also lines up with the jumps in the error at the same time steps we have seen in Fig. 14. The fact that the runtime per step decreases with each step up to the next bootstrapping operation also matches the performance characteristics we saw earlier in Sec. 3, where many FHE operations became slower with increasing multiplicative depth. This is due to the ciphertext modulus being reduced after each multiplication, which in turn reduces the number of computational steps required for each subsequent FHE operation. The same behavior can be observed here, where the remaining multiplicative depth decreases with each time step until it is refreshed by bootstrapping. A similar trend can be observed in Fig. 15(b) for the 2D simulations.
(a) 1D.
(b) 2D.
Figure 15: Execution times of secure numerical simulations.
To highlight their interdependence, in Fig. 16 we analyze the error and runtime of the secure simulations with the Lax-Wendroff scheme in 2D for different multiplicative depths after bootstrapping $l_{\text{refresh}}$. The error becomes smaller with larger $l_{\text{refresh}}$, aligning with our previous observation that the bootstrapping incurs a high approximation error. The overall runtime first decreases with increasing multiplicative depth, which is expected since the bootstrapping operation is less frequent. However, there seems to exist an optimum value, since after approximately $l_{\text{refresh}}=17$ the runtime increases again. This confirms our earlier discussion that there exists a trade-off between the computational cost of the FHE operations, which increases with larger values for $l_{\text{refresh}}$, and the frequency of the bootstrapping operation. The optimal multiplicative depth thus depends on the specific application and the used numerical scheme.
Figure 16: Error and execution time for secure simulations with different values of $l_{\text{refresh}}$ (multiplicative depths after bootstrapping). The simulations of the 2D linear advection equation were performed with the Lax-Wendroff scheme.
Finally, we examine the effect of using the iterative bootstrapping technique with two iterations as discussed in Sec. 3.7. Compared to the previous results with standard bootstrapping in Fig. 14, Fig. 17 shows that the error is reduced by three orders of magnitude, from approximately $\mathcal{O}(10^{-6})$ to $\mathcal{O}(10^{-9})$. This matches our previous findings in Fig. 8, reemphasizing the dominant role of the bootstrapping operation on the overall error caused by the CKKS scheme. At the same time, the runtime per step rises from $20.8$ s to $30.4$ s. The increase in runtime is mainly due to doubling the number of bootstrapping operations. Furthermore, for iterative bootstrapping, a larger number of available levels $l=2$ is required (compared to $l=1$ before, see also Sec. 3.7). Nevertheless, it is clear that iterative bootstrapping can be a viable option in scenarios where the error introduced by the CKKS scheme is a limiting factor and the computational cost is not prohibitive. Interestingly, the previously observed self-healing capabilities of the finite difference scheme seem to disappear with iterative bootstrapping, as the error no longer decreases between two bootstrapping operations.
Figure 17: Error (left) and execution time (right) for secure simulations with iterative bootstrapping. The simulations of the 2D linear advection equation were performed with the Lax-Wendroff scheme.
Overall, our findings demonstrate that it is possible to implement secure numerical simulations with the CKKS scheme, and that we are able to obtain sufficiently accurate results with the implemented numerical methods for the linear advection equation. They also show that the CKKS scheme is a significant source of errors and that it introduces a considerable performance overhead. Especially the bootstrapping operation has a large impact on both error and runtime. For practical applications, it thus seems advisable to primarily focus on optimizing the multiplicative depth and the bootstrapping configuration to minimize the introduced error and to maximize the performance of secure simulations. In scenarios where runtime is not the primary limitation, iterative bootstrapping may be used to improve the accuracy of FHE-based simulations.
5.4 Parallel computations
As evident from the previous sections, FHE operations incur a considerable performance overhead. Therefore, it makes sense to evaluate potential ways to speed up the secure computations. One possible approach is to parallelize the FHE operations.
The OpenFHE library internally uses the OpenMP library for multithreading computations, which is also available in Julia. To study the efficiency of CKKS parallelization, we consider the most expensive numerical solver from this work, i.e., the Lax-Wendroff scheme for the linear advection equation in 2D, and execute it for various numbers of OpenMP threads. This can be controlled by setting the environment variable OMP_NUM_THREADS to the appropriate number of threads before starting the Julia runtime.
Fig. 18 shows the runtimes for the parallel execution of the secure simulations. For the used CKKS parameters and hardware setup, we could achieve at most a three-fold speedup in computation time with eight threads, and a ten-fold speedup with $32$ threads in configuration time. In either case, using more threads than this did not lead to any significant improvements in runtime. Since the multithreading capabilities of the OpenFHE library are readily available and easy to use, it is advisable to employ them to speed up the CKKS computations whenever possible.
(a) Execution time for different numbers of OpenMP threads.
(b) Speedup for different numbers of OpenMP threads.
Figure 18: Parallel execution with multiple OpenMP threads of the Lax-Wendroff scheme for the linear scalar advection equation in 2D. Runtime and speedup are analyzed separately for the configuration and the computation phase.
6 Extension to other mathematical models and numerical methods
As demonstrated in the previous section, secure numerical simulations are generally possible. It was also noted that the primary impediments to the practical use of the CKKS scheme for such scenarios are the large computational overhead and, to a lesser extent, the precision loss. Apart from this, three main challenges arise when considering to extend the simulation prototypes shown in this paper to other PDEs and to more complex numerical schemes: a single built-in data structure, limited function evaluation capabilities, and constraints in control flow logic when it comes to conditional branching. In the following, we will discuss these issues and potential solutions to them, to give other researchers a starting point for their own investigations into secure numerical simulations. We will conclude this section with a brief summary of some general strategies for improving performance.
6.1 Data structures
A vector of real numbers is the only encrypted data structure available in the CKKS scheme, as described in Sec. 2.1. More complex data structures, such as tensors or matrices, are not natively supported. In Sec. 4.3.2, we have introduced one potential approach to a matrix representation for use with explicit, mesh-based numerical simulations. In other algorithms that involve matrix-matrix or matrix-vector multiplications, specialized encoding strategies for the matrix lead to more efficient algorithms in terms of the multiplicative depth [55, 56, 28].
In general, the concrete choice of data representation depends on the target application and the used numerical method. Due to the significant computational overhead of the CKKS scheme, it is necessary to make full use of its batch processing capabilities, i.e., the ability to do element-wise addition and multiplication on the entire data vector. Therefore, discretization schemes that allow for efficient vectorization are preferable when considering other numerical methods for secure simulations. The choice of PDE is largely unaffected by the data structure, as long as a single model with a constant number of state variables is used throughout the computational domain.
6.2 Function evaluation
As discussed before, the CKKS scheme supports only a limited set of arithmetic operations, i.e., element-wise addition and multiplication of the ciphertexts, as well as rotation. This means that more complex arithmetic operations, such as division, exponentiation, or trigonometric functions, are not directly available but must be approximated by a series of simpler operations or using a polynomial interpolation. For example, the division of two ciphertexts can be implemented with the Newton-Raphson method [35] or Goldschmidt’s algorithm [36], the latter of which is also used to implement floating point division in some CPU families [57]. Exponentiation or trigonometric functions may be approximated by a Taylor series expansion [58].
Since evaluating a polynomial expression is straightforward with the built-in arithmetic operations, any function that is reasonably well approximated by a polynomial can be computed with sufficient accuracy. For many other functions, an iterative approach is required. Here, the main difficulty is that the number of iterations must be fixed a priori, since branching based on encrypted data is not possible (refer to the next section for details). Without fundamental limitations to computing branch-free functions with FHE, the primary challenge is therefore to find an efficient approximation that converges robustly. Selecting the best approach is problem-dependent and likely requires some experimentation to find a good balance between accuracy and performance. When considering other mathematical models or numerical discretizations, it is thus advisable to choose options that minimize the multiplicative depth of the FHE operations to avoid costly bootstrapping operations.
For many discontinuous functions, e.g., sign/comparison, the use of CKKS is challenging as the polynomial approximation in the proximity of the singularity point(s) requires extremely high-degree polynomials, i.e., very large multiplicative depth and runtime. Instead, other FHE schemes are often used, such as DM [23] or CGGI [24]. DM and CGGI support the evaluation of arbitrary functions via lookup tables using a technique called functional or programmable bootstrapping [59]. In these scenarios, a scheme switching procedure is employed [60, 61], where the encryption of the data is temporarily changed from CKKS to DM/CGGI. This allows for the evaluation of the discontinuous function, after which the data is converted back to a CKKS representation. Note that such scheme switching capability is supported in OpenFHE [40]. The main drawback of scheme switching is the significant computational overhead, especially when high precision is required [62, 51, 63]. A promising new method for evaluating discontinuous functions using functional CKKS bootstrapping is proposed and implemented in [64]. The method enables arbitrary function evaluation via lookup tables and achieves a throughput that is orders of magnitude higher than DM or CGGI.
6.3 Conditional branching
By their very nature, many of the control flow directives commonly used in regular programs cannot be employed with FHE methods. More specifically, FHE precludes the use of conditional statements or branches—including loops with a termination condition—that depend on encrypted data. This is due to the fact that the encryption of the data prevents the processor from knowing the actual values of the data. If this were not the case, it would be possible for a malicious actor to infer information about the encrypted data, e.g., by using a bisection algorithm to compare the secret ciphertext with known values.
As noted in [65], this restriction in control flow logic presents a fundamental challenge when converting regular programs to FHE algorithms. Since branching on an encrypted value is impossible by design, an FHE algorithm encountering a conditional branch must therefore first compute the outcome for each possible branch. The final result is then constructed by combining the results from all branches and applying an appropriate mask that selects the correct result. For example, Alg. 4 shows a simple branching statement in a regular program. Based on a function condition() that returns either $1$ or 00, the value of $r$ is computed by calling either foo() or bar(). The equivalent FHE algorithm in Alg. 5 first computes the value of the condition, evaluates both branches, and then combines the results with the condition value to obtain the final result.
Algorithm 4 Regular conditional branching.
[…]
1:$r\leftarrow\text{foo}()$Else
2:$r\leftarrow\text{bar}()$EndIf
[…]
Algorithm 5 Conditional branching with FHE.
1:$c\leftarrow\text{condition()}$()
2:$r_{\text{foo}}\leftarrow\text{foo}()$
3:$r_{\text{bar}}\leftarrow\text{bar}()$
4:$r\leftarrow c*r_{\text{foo}}+(1-c)*r_{\text{bar}}$
Such an approach is feasible if a small, predetermined number of branches needs to be considered, e.g., when computing the well-known HLLC flux for the compressible Euler equations, which has different branches based on three different wave speeds [66]. It becomes intractable, however, if the number of branches is large and possibly unknown in advance, e.g., when terminating a time step loop while using a dynamic, CFL-based time step size. Furthermore, numerical schemes that strictly require conditional branching based on current solution values are not directly compatible with FHE. This includes methods that dynamically modify their data structures during the simulation, such as periodic remeshing when dealing with deforming boundaries or adaptive mesh refinement. As a result, this poses a concrete limitation on the applicability of FHE to certain classes of numerical simulations.
Therefore, when choosing a numerical method for secure numerical simulations, it is important to avoid the need for control flow that depends on encrypted data and that does not have a fixed number of branches. While the statement that FHE can be used to compute any function (see Sec. 2) is not violated, it means that in practice, the translation to an FHE algorithm is not always feasible. As a workaround, one can consider allowing certain control flow operations to be unencrypted, like we did in Sec. 4.4 for the time loop in Alg. 3.
6.4 General strategies for improving performance
As noted in the previous subsections, most methods for dealing with mathematical models and numerical schemes that require more complex FHE operations incur a significant computational overhead. To mitigate this, several strategies can be employed to optimize the performance of FHE algorithms:
  • 1.
    Prefer parallelizable algorithms over sequential ones. As discussed in Sec. 2.1.6, using a binary approach for multiple multiplications can reduce the overall multiplicative depth of the FHE operations. This naturally applies also to all other functions that consume multiplicative levels. For example, consider loop unrolling to manually encode its iterations more efficiently.
  • 2.
    Minimize or avoid FHE-hard operations. For instance, another set of state variables, e.g., using a different non-dimensionalization strategy, may save some (costly) division operation in each time step.
  • 3.
    Reduce the number of rotations. While they do not consume multiplicative levels, they require similar computational resources as multiplications (see Table 2). For example, summation over a row/column of a matrix may be expressed as a series of additions instead.
  • 4.
    Consider performing costly computations a priori. For example, it might be more efficient to use a fixed, smaller time step based on an conservative estimate, rather than computing a potentially larger step size dynamically in each time step loop iteration.
A more detailed description of these and other optimization strategies can be found in [28].
7 Conclusions and outlook
In this work, we presented the first prototype of secure numerical simulations using fully homomorphic encryption. We gave an overview of the CKKS scheme and its efficient implementation within the OpenFHE library, and introduced our OpenFHE.jl and SecureArithmetic.jl packages, which provide a user-friendly interface to OpenFHE in the Julia programming language. A detailed analysis of the accuracy and performance of the individual CKKS routines revealed that while each operation introduces some numerical error, bootstrapping is by far the least accurate procedure. Similar observations were made for the runtime behavior, where the bootstrapping mechanism is orders of magnitude slower than other CKKS operations.
Next, we implemented the first-order upwind and second-order Lax-Wendroff schemes for the linear advection equations for secure numerical simulations using the SecureArithmetic.jl package. For this purpose, we introduced a $\operatorname{circshift}$ operation that is compatible with the CKKS scheme and that works for both 1D and 2D data representations. Convergence tests verified the correctness of our implementations and showed that the CKKS encryption does not break the convergence properties of the finite difference schemes.
We then conducted homomorphically encrypted numerical simulations by building upon the tools and methods developed in the previous sections. Our results illustrate that it is feasible to implement secure numerical simulations with the CKKS scheme, with results that match the unencrypted simulations well. However, the encrypted operations introduce a large computational overhead and are a significant source of error, to the point that we were able to achieve less than single-precision floating point accuracy at the cost of a six-orders-of-magnitude increase in runtime. Nearly all of these undesirable effects can be attributed to the bootstrapping procedure, which is the most inaccurate and slowest operation in the CKKS scheme. However, bootstrapping is also necessary to refresh the multiplicative depth of the ciphertexts, making it indispensable for long-running simulations. Moreover, iterative bootstrapping can further improve the accuracy of an FHE algorithm, though at a non-negligible runtime penalty that needs to be taken into account. We also showed that the CKKS scheme implementation in OpenFHE is parallelizable, and that using multiple threads can lead to a significant reduction in the solution time.
Finally, we explored challenges—and how to address them—when extending secure numerical simulations beyond the presented examples. Many of the discussed issues can be mitigated through careful algorithm design and by accepting some computational overhead. However, the lack of dynamic, value-based conditional branching precludes the use of schemes that strictly depend on it, such as adaptive mesh refinement.
To summarize, our work demonstrates that it is possible to perform secure numerical simulations using fully homomorphic encryption. At the same time, the large computational overhead and precision loss in the bootstrapping operation are significant obstacles that need to be taken into account. Therefore, currently FHE-secured numerical simulations only make practical sense in specialized applications where the security of the data is of utmost importance. Thus, more detailed investigations are necessary to optimize the use of the CKKS scheme for numerical simulations. By providing a user-friendly interface to the OpenFHE library in Julia, and by implementing some important building blocks such as the $\operatorname{circshift}$ algorithms, we hope to lay the foundation for further research in this direction. In addition, our discussion of the challenges and potential solutions for extending the secure numerical simulations to other mathematical models and numerical methods should provide a starting point for other researchers interested in this field.
In the future, we plan to extend our work in several areas. We would like to increase the complexity of the simulations by looking at systems of equations and at three-dimensional problems. This will also necessitate an extension of the $\operatorname{circshift}$ operation to support multiple ciphertexts and tensors. Furthermore, we will investigate the discretization of nonlinear equations, which will require us to handle additional arithmetic operations such as division and exponentiation. Moreover, we would like to better understand the observed behavior that the accuracy of secure numerical simulations with bootstrapping can improve over time. Finally, we plan to continue developing the SecureArithmetic.jl package as a tool for experimentation and rapid prototyping with encrypted algorithms, to open up FHE to a wider audience in the computational science community.
Data availability
All code used in this work and the data generated during this study are available in our reproducibility repository [30].
CRediT authorship contribution statement
Arseniy Kholod: Data curation, Investigation, Methodology, Software, Validation, Visualization, Writing –- original draft, Writing — review & editing. Yuriy Polyakov: Methodology, Validation, Writing — review & editing. Michael Schlottke-Lakemper: Conceptualization, Investigation, Methodology, Resources, Software, Supervision, Validation, Writing –- original draft, Writing — review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to enhance readability. They further used the GitHub Copilot plugin for Visual Studio Code for assistance with mathematical notation and formatting in LaTeX. After using these tools/services, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Acknowledgments
We gratefully acknowledge the computing time provided by Numerical Simulation research group of Gregor Gassner, University of Cologne. Michael Schlottke-Lakemper further acknowledges funding through the DFG research unit FOR 5409 "Structure-Preserving Numerical Methods for Bulk- and Interface Coupling of Heterogeneous Models (SNuBIC)" (project number 463312734), as well as through a DFG individual grant (project number 528753982). Moreover, we would like to thank Andrew Hazel for his valuable feedback on the manuscript.